log6(b2 + 2) + log62 = 2


Answer by mukhopadhyay(490)

log6(b^2 + 2) + log62 = 2
=>log(base6)2(b^2+2) = 2
=>2(b^2+2)=36
=> b^2+2 = 18
=> b^2 = 16
=> b=4 or b=-4
Answer: b=4 or b=-4